OpenStudy (anonymous):

PLEASE HELP Which ordered pair is the solution to the system of equations? x+y=4 x-y=-6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4) @ganeshie8 @hartnn @mathstudent55 @Abhisar @dan815 @iambatman @Callisto @.Sam.

OpenStudy (anonymous):

x+y = 4 x - y = - 6 subtract second eq from first (x+y) -(x-y) = 4-(-6) 2y = 10 y = 5 Put y = 5 in any equation to find value of x

Miracrown (miracrown):

Since you have answers, you could plug in each (x, y) pair to see which makes both equations true. If you want to solve it without plugging in, you would use either elimination or substitution to solve the system

OpenStudy (anonymous):

is it b @Miracrown @sangya21

OpenStudy (anonymous):

which coordinate have y = 5?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

omg thank you so much could u help me with 4 more ??@sangya21

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-y=-5 x+3y=22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

@sangya21